OpenStudy (superhelp101):

A solution is made by dissolving 1.00 moles of sodium chloride (NaCl) in 155 grams of water. If the molal boiling point constant for water (Kb) is 0.51 °C/m, what would be the boiling point of this solution? Show all of the work needed to solve this problem.

OpenStudy (superhelp101):

@Abmon98

OpenStudy (superhelp101):

@thomaster @paki

OpenStudy (joannablackwelder):

ΔT = Kb*m

OpenStudy (joannablackwelder):

Can you find the molality of the solution?

OpenStudy (superhelp101):

molality = 1.00 mol/ 0.155 Kg = 6.45 delta T = m x kb x i i = Vant'Hoff factor for NaCl = 2 delta T = 6.45 x 0.51 x 2 = 6.6 boiling point = 100 + 6.1 = 106.1 °C this is what i did is it right?

OpenStudy (joannablackwelder):

Yes, I get the delta T to be 6.6. 100+6.6 is what?

OpenStudy (superhelp101):

106.6?

OpenStudy (joannablackwelder):

:)

OpenStudy (superhelp101):

oh so is the boiling point is is 106.6?

OpenStudy (joannablackwelder):

That's what I get.

OpenStudy (superhelp101):

thank you!!

OpenStudy (joannablackwelder):

No worries. :)